OpenStudy (anonymous):

Which ordered pair is a solution of y = x - 4.

OpenStudy (anonymous):

what are the ordered pairs

OpenStudy (anonymous):

0,-4 5,-1 3,-7 -7,-3

OpenStudy (anonymous):

y=x-4 0,-4 -4=0-4 5,-1 -1=5-4 3,-7 -7=3-4 -7,-3 -3=-7-4 now just solve the left sides of the equation and what ever one works is the answer

OpenStudy (anonymous):

thank you .